I submitted a similar problem earlier, but it had a typo and could not be understood. I'm going to retype the problem with hopefully no typos this time. Could you please help me solve this quadratic equation using completing the square method? x^2+5x+6=0
Answer by Alan3354(69443)

solve this quadratic equation using completing the square method?
x^2+5x+6=0
-----------
x^2+5x = -6
a = 1, b = 5
Add (b/2a)^2 to each side
x^2 + 5x + (5/2)^2 = -6 + (5/2)^2
x^2 + 5x + 25/4 = -6 + 25/4 = 1/4
(x + 5/2)^2 = (1/2)^2
x + 5/2 = 1/2
x = -2
================
x + 5/2 = -1/2
x = -3